OpenStudy (anonymous):

-5(1 - 5x) + 5(-8x - 2) = -4x - 8x I got x = -5, but I don't feel like it's correct. Help?

jimthompson5910 (jim_thompson5910):

x = -5 is the correct answer

jimthompson5910 (jim_thompson5910):

Check: -5(1 - 5x) + 5(-8x - 2) = -4x - 8x -5(1 - 5(-5)) + 5(-8(-5) - 2) = -4(-5) - 8(-5) -5(1 +25) + 5(40 - 2) = 20 + 40 -5(26) + 5(38) = 20 + 40 -130 + 190 = 20 + 40 60 = 60 So the answer checks out

OpenStudy (anonymous):

Oh wow I really need to pay more attention to what I'm doing... thanks again

jimthompson5910 (jim_thompson5910):

yw, you're getting the right answer, which is great

jimthompson5910 (jim_thompson5910):

you just have to get used to checking it and trusting yourself